OpenStudy (anonymous):

help pls! with this question -> with an average acceleration of -0.50m/s^2 to reach a speed of 12.0m/s. how long does it take for this acceleration to occur?

OpenStudy (anonymous):

\[v=v_{o}+a \times t\]

OpenStudy (anonymous):

12 = 0 - 0.5*t

OpenStudy (anonymous):

24 seconds

OpenStudy (anonymous):

Since it's -0.5 m/s^2 it's deceleration

OpenStudy (anonymous):

If the object was going at 12 m/s to begin with, and it started decelerating at -0.5 m/s^2 - it would also take it 24 seconds to stop.

OpenStudy (anonymous):

how do u get 24 ?

OpenStudy (anonymous):

12 = 0.5*t /: 0.5 24 = t

OpenStudy (anonymous):

With what speed does the object start? 0? not 0?

OpenStudy (anonymous):

is -0.5

OpenStudy (anonymous):

If it starts at rest, it's an illogical question.

OpenStudy (anonymous):

mmmm still not getting it lol i suck on math

OpenStudy (anonymous):

thanks anyways :)

OpenStudy (anonymous):

Do you have the full question? or is that all they give you?

OpenStudy (anonymous):

is all that they give me

OpenStudy (anonymous):

lemme see if a get it so i have to subtract 0-(-0.5) divde by 12 right ?

OpenStudy (anonymous):

according to the formula it's add, so the time will result in a negative number.

OpenStudy (anonymous):

mkay thank you.